# Gradient Coupling: The Hidden Barrier to Generalization in Agentic Reinforcement Learning

### v3

@@ 1 Introduction
Currently, with the rapid development of large language models (GPT-4; gemini; qwen; deepseek_r1; instruction_following), building an autonomous agent capable of solving real world, complex and long-horizon tasks has gained significant attention given the great power of large language models (agent_first; agent_smart; world_device_control_agent; CodeAgent; empathetic_agents; gui_agent). However supervised fine-tuning shows unsatisfying performance and can easily cause catastrophic forgetting (RL_generalize) when faced with complex agentic tasks. Therefore, reinforcement learning is now the most effective method to solve multi-turn interaction especially long horizon tasks which require lots of steps to solve (GiGPO; agentevolver; multi_agent_rl).
Outcome-based reinforcement learning methods, such as GRPO (GRPO; DAPO; reinforce++; logic_rl), have shown promising performance in domains like mathematical reasoning. Yet, we find that, these methods underperform in multi-turn interactive agentic tasks. Reinforcement Learning can greatly improve the performance on trained environments but the agent tends to perform poorly on unseen environments. If we modify the trained environment, then the model might completely fail to do the job and result in unreasonable actions. For example, an agent is trained on tasks requiring it to pick up a single object and place it on a table. During testing, if the task is modified to "pick up two objects and place them on the table," the agent—over-reliant on its learned policy—might pick up only the first object and then terminate the task. This action, while optimal within the context of its training data, constitutes a failure in the new, more complex test scenario. This example highlights how policies that are effective in a narrow training setting can be brittle and fail to adapt to novel task requirements.
@@ 2 Related Work
Reinforcement Learning for LLM Reinforcement Learning algorithms like PPO (PPO) is growing extremely popular because it can greatly improve the performance (instruction_following). Then Direct Preference Optimization (DPO) (DPO) is proposed to simplify the optimization process of PPO, and GRPO removes the critic in PPO and greatly enhances the reasoning performance of LLM by estimating advantages by using batches of samples generated from the same prompt which is also widely applied in tasks like mathematical reasoning (GRPO), retrieval (atom_searcher; do_not_abstain; search-R1) and tool use (toolRL; CodeAgent) and multi-turn agent tasks (RAGEN; long_horizon_RL_apple; RLVMR; harnessing_uncertainty; sweetRL). However, we find that although reinforcement learning can greatly help the agent to perform good in the trained task, it might fail to generalize to similar but unseen tasks.
Agent Generalization The agent trained with RL shows great performance on in domain tasks, but the generalization is poor. SFT_memorize_RL_generalize claims that RL shows better generalization than SFT, but rl_no_gene argue that RL fails to improve the generalization of LLM. early_experimence use self reflection and future prediction to improve the generalization. While agentevolver automatically collect online data and collect data to further fine-tune the model, improving the performance. And RLVMR try to design meta thinking prompts to improve the generalization. And some researchers try to use multi agent interaction to improve the generalization (multi_agent_rl; rema_multi_agent) But they did not reveal why current methods fail to generalize, only try to solve it through prompt design and data collection while our method try to reveal why agents fail to generalize and solve it.
@@ 4 Methodology
Our analysis shows that gradient coupling could have negatively interference on agentic RL. The core objective, therefore, is to decouple the learning signals for positive and negative trajectories, ensuring that correct actions are reinforced while incorrect ones are suppressed. While prior work like GRPO_squeeze has attempted to mitigate such interference, its methodology is ill-suited for agentic tasks and they does not consider that the gradient coupling can somehow help generalization by connecting positive samples. They try to downweight penalties on specific "harmful" tokens within negative samples to reduce gradient coupling, this strategy is effective in domains like mathematical reasoning, where harmful similarity often originates from generic logical transition words (e.g., "so," "then"). Penalizing these tokens has minimal impact on the core reasoning process.
@@
As shown in GRPO_squeeze, the gradient coupling in GRPO can be expressed as:
@@
Reward Allocation: Each generated judgment $j_{i}^{(k)}$ is assigned a reward, $r_{\text{GCD}}^{(k)}$. This reward is simply 1 if the sentiment of the judgment $j_{i}^{(k)}$ matches the ground-truth label $y_{i}$, and 0 otherwise. About how to obtain the label of good/bad actions. Inspired by GiGPO, if the current observation reoccurs in subsequent steps, this suggests that the action might be incorrect, as it leads to a previously visited state, potentially indicating a loop or redundant move. Otherwise if the current step is the last one, it means that it directly leads to success, so we recognize it as a correct action. To improve data diversity, we also use DeepSeek V3 to generate labels on actions. And we only select those label with high confidence, specifically, we prompt the model for 5 times, if it answer consistently for 4 or more times, then it is used as a training sample. Then we prompt the actor model to judge if the action is good one and allocate reward to conduct GRPO training.
@@ 5 Experiments
We conduct experiments on ALFWorld (alfworld) and ScienceWorld (agent_smart), and we train Qwen2.5-3B-Instruct and Qwen2.5-7B-Instruct on the environments fro 150 epochs. We update the suggestion every 10 steps during RL, and we do not add the generated suggestion when testing for fair comparison. We only add 30 samples to conduct judgment, so it does not affect the efficiency, and we discuss the time consumption and the performance of the judgment task on in/out domain tasks in Appendix A.
We present the results in Table 2, the result shows that GCD can help the performance of both in domain tasks and out domain tasks. And we can also observe that NTHR (GRPO_squeeze) who try to downweight penalties on specific "harmful" tokens within negative samples fails to help. As we show in Figure 7, the relationship between different samples is caused by lots of meaningful reasoning or actions tokens, penalizing these tokens would hinder the effect of negative gradients, leading to suboptimal performance.
@@ 5.2 Difference with Turn Level Reward
At first glance, our method’s use of turn-level judgments appears related to recent work on turn-level rewards, such as GiGPO (GiGPO) and RLVMR (RLVMR), which also evaluates intermediate actions. However, we contend that our approach targets a fundamentally different problem. These prior methods primarily address the challenge of sparse rewards and credit assignment. Their goal is to propagate learning signals to crucial intermediate steps within a trajectory, for example, by rewarding beneficial actions even if the overall outcome was a failure.
@@ A.1 Datasets
We conduct experiments on ALFWorld (alfworld) and ScienceWorld (scienceworld), ALFWorld is a household environment which requires the agent to explore the room and accomplish tasks like "pick up a pencil and put it on the table". ScienceWorld is designed to evaluate the scientific reasoning abilities with 10 scientific domains and 30 subcategories. Regarding the in-domain and out-of-domain split, following RLVMR, for ALFWorld, we use Cool & Place and Pick Two & Place as the out domain task and other 4 tasks as in domain task; for ScienceWorld, the final task type of each topic is reserved as out domain task evaluation.

### v4

@@ 1 Introduction
Currently, with the rapid development of large language models (Achiam et al., 2023; Team et al., 2024; Bai et al., 2023; Guo et al., 2025; Ouyang et al., 2022), building an autonomous agent capable of solving real world, complex and long-horizon tasks has gained significant attention given the great power of large language models (Zeng et al., 2023; Wang et al., 2022a; Bai et al., 2024; Zhang et al., 2024; Wang et al., 2025b, 2024). However supervised fine-tuning shows unsatisfying performance and can easily cause catastrophic forgetting (Fu et al., 2025) when faced with complex agentic tasks. Therefore, reinforcement learning is now the most effective method to solve multi-turn interaction especially long horizon tasks which require lots of steps to solve (Feng et al., 2025; Zhai et al., 2025; Liao et al., 2025).
Outcome-based reinforcement learning methods, such as GRPO (Shao et al., 2024; Yu et al., 2025; Hu et al., 2025; Xie et al., 2025), have shown promising performance in domains like mathematical reasoning. Yet, we find that, these methods underperform in multi-turn interactive agentic tasks. Reinforcement Learning can greatly improve the performance on trained environments but the agent tends to perform poorly on unseen environments. If we modify the trained environment, then the model might completely fail to do the job and result in unreasonable actions. For example, an agent is trained on tasks requiring it to pick up a single object and place it on a table. During testing, if the task is modified to "pick up two objects and place them on the table," the agent—over-reliant on its learned policy—might pick up only the first object and then terminate the task. This action, while optimal within the context of its training data, constitutes a failure in the new, more complex test scenario. This example highlights how policies that are effective in a narrow training setting can be brittle and fail to adapt to novel task requirements.
@@ 2 Related Work
Reinforcement Learning for LLM Reinforcement Learning algorithms like PPO (Schulman et al., 2017) is growing extremely popular because it can greatly improve the performance (Ouyang et al., 2022). Then Direct Preference Optimization (DPO) (Rafailov et al., 2023) is proposed to simplify the optimization process of PPO, and GRPO removes the critic in PPO and greatly enhances the reasoning performance of LLM by estimating advantages by using batches of samples generated from the same prompt which is also widely applied in tasks like mathematical reasoning (Shao et al., 2024), retrieval (Deng et al., 2025b; Liu et al., 2025; Jin et al., 2025a) and tool use (Qian et al., 2025; Zhang et al., 2024) and multi-turn agent tasks (Wang et al., 2025c; Chen et al., 2025; Zhang et al., 2025b; Wang et al., 2025a; Zhou et al., 2025). However, we find that although reinforcement learning can greatly help the agent to perform good in the trained task, it might fail to generalize to similar but unseen tasks.
Agent Generalization The agent trained with RL shows great performance on in domain tasks, but the generalization is poor. Chu et al. (2025) claims that RL shows better generalization than SFT, but Jin et al. (2025b) argue that RL fails to improve the generalization of LLM. Zhang et al. (2025a) use self reflection and future prediction to improve the generalization. While Zhai et al. (2025) automatically collect online data and collect data to further fine-tune the model, improving the performance. And Zhang et al. (2025b) try to design meta thinking prompts to improve the generalization. And some researchers try to use multi agent interaction to improve the generalization (Liao et al., 2025; Wan et al., 2025) But they did not reveal why current methods fail to generalize, only try to solve it through prompt design and data collection while our method try to reveal why agents fail to generalize and solve it.
@@ 4 Methodology
Our analysis shows that gradient coupling could have negatively interference on agentic RL. The core objective, therefore, is to decouple the learning signals for positive and negative trajectories, ensuring that correct actions are reinforced while incorrect ones are suppressed. While prior work like Deng et al. (2025a) has attempted to mitigate such interference, its methodology is ill-suited for agentic tasks and they does not consider that the gradient coupling can somehow help generalization by connecting positive samples. They try to downweight penalties on specific "harmful" tokens within negative samples to reduce gradient coupling, this strategy is effective in domains like mathematical reasoning, where harmful similarity often originates from generic logical transition words (e.g., "so," "then"). Penalizing these tokens has minimal impact on the core reasoning process.
@@
As shown in Deng et al. (2025a), the gradient coupling in GRPO can be expressed as:
@@
Reward Allocation: Each generated judgment $j_{i}^{(k)}$ is assigned a reward, $r_{\text{GCD}}^{(k)}$. This reward is simply 1 if the sentiment of the judgment $j_{i}^{(k)}$ matches the ground-truth label $y_{i}$, and 0 otherwise. About how to obtain the label of good/bad actions. Inspired by Feng et al. (2025), if the current observation reoccurs in subsequent steps, this suggests that the action might be incorrect, as it leads to a previously visited state, potentially indicating a loop or redundant move. Otherwise if the current step is the last one, it means that it directly leads to success, so we recognize it as a correct action. To improve data diversity, we also use DeepSeek V3 to generate labels on actions. And we only select those label with high confidence, specifically, we prompt the model for 5 times, if it answer consistently for 4 or more times, then it is used as a training sample. Then we prompt the actor model to judge if the action is good one and allocate reward to conduct GRPO training.
@@ 5 Experiments
We conduct experiments on ALFWorld (Shridhar et al., 2020) and ScienceWorld (Wang et al., 2022a), and we train Qwen2.5-3B-Instruct and Qwen2.5-7B-Instruct on the environments fro 150 epochs. We update the suggestion every 10 steps during RL, and we do not add the generated suggestion when testing for fair comparison. We only add 30 samples to conduct judgment, so it does not affect the efficiency, and we discuss the time consumption and the performance of the judgment task on in/out domain tasks in Appendix A.
We present the results in Table 2, the result shows that GCD can help the performance of both in domain tasks and out domain tasks. And we can also observe that NTHR (Deng et al., 2025a) who try to downweight penalties on specific "harmful" tokens within negative samples fails to help. As we show in Figure 7, the relationship between different samples is caused by lots of meaningful reasoning or actions tokens, penalizing these tokens would hinder the effect of negative gradients, leading to suboptimal performance.
@@ 5.2 Difference with Turn Level Reward
At first glance, our method’s use of turn-level judgments appears related to recent work on turn-level rewards, such as GiGPO (Feng et al., 2025) and RLVMR (Zhang et al., 2025b), which also evaluates intermediate actions. However, we contend that our approach targets a fundamentally different problem. These prior methods primarily address the challenge of sparse rewards and credit assignment. Their goal is to propagate learning signals to crucial intermediate steps within a trajectory, for example, by rewarding beneficial actions even if the overall outcome was a failure.
@@ A.1 Datasets
We conduct experiments on ALFWorld (Shridhar et al., 2020) and ScienceWorld (Wang et al., 2022b), ALFWorld is a household environment which requires the agent to explore the room and accomplish tasks like "pick up a pencil and put it on the table". ScienceWorld is designed to evaluate the scientific reasoning abilities with 10 scientific domains and 30 subcategories. Regarding the in-domain and out-of-domain split, following Zhang et al. (2025b), for ALFWorld, we use Cool & Place and Pick Two & Place as the out domain task and other 4 tasks as in domain task; for ScienceWorld, the final task type of each topic is reserved as out domain task evaluation.